OpenStudy (anonymous):

What is the solution to the following system of equations? 2x + 3y = 6 x + y = 4 The solution of the system of equations is A1_______, A2______?

OpenStudy (anonymous):

2x+3y=6 is the first equation x+y =4 is the second equation

OpenStudy (anonymous):

multiply by two for the seconf equation

OpenStudy (anonymous):

u will get 2x+2y=8

OpenStudy (anonymous):

this will be the 3rd equation

OpenStudy (anonymous):

so finally u take the 1st equation to minus the 3rd equation

OpenStudy (anonymous):

2x+3y-(2x=2y)=6-8

OpenStudy (anonymous):

sory , 2x+3y-(2x+2y)=6-8

OpenStudy (anonymous):

so y=-2 , sub this into the 2nd equation and u will get x =6 , solve